-1 points CJ10 20 P088 My Notee Ask Y O Ask The col of wire in a galvanom eter has a resistance of RC-73.0 Ω. The galvanometer exhibits a fut scale defection when the eument through it is 0270 mA A resistor is connected in ser combination so as to produce a voltmeter. The voit ter is to have a full-scale deflection when it measures a potential difference of 13.0 V. what is the resistance of this resistor? Addsional Materials secton 20 11
Answer #1

given Rc = 73 Ohm

Ig = 0.270 mA = 0.27 X 10-3 A

V = 13 volts

R = ?

using V = Ig ( R + Rc )

13 = 0.27 X 10-3 X ( R +73 )

( R +73 ) = 13 / 0.27 X 10-3

( R +73 ) = 48148.148

R = 48075.148 Ohm